what the answer of these using elimination method
3x-3000y=3
2x+1y=1100

Answer by TimothyLamb(4379)

3x-3000y=3
2x+1y=1100
---
put the system of linear equations into standard form
---
3x - 3000y = 3
2x + y = 1100
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 549.72564
y = 0.54872564
---
by elimination (of the y term):
---
3x - 3000y = 3
(2x + y = 1100)*3000
---
3x - 3000y = 3
6000x + 3000y = 1100*3000
--
add the two equations:
---
6003x = 1100*3000 + 3
x = (1100*3000 + 3)/6003
x = 549.72564
---
3*(549.72564) - 3000y = 3
3000y = 3(549.72564) - 3
y = (3*(549.72564) - 3)/3000
y = 0.54872564
---
solution:
x = 549.72564
y = 0.54872564
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php